Following electrically operated equipments are running as given below. Calculate the energy bill for the month of March, if the energy charges are Rs 525/- per unit Fluorescent Tube 40 Watt, 8 Hrs/day

Fan- 60 Watt, 10 Hrs/day

Geyser -1000 Watt, I Hr/day

Ac-1500 Watt, 4 Hrs/day
1
Expert's answer
2021-07-12T03:36:54-0400

"Energy = Power*time\\"

Energy consumed by Fluorescent Tube 40 W per day "= 8\u221740 Wh \n\n = 0.32kWh"

Energy consumed by Fan 60 W per day "= 8\u221760 Wh\n\n = 0.48 kWh"

Energy consumed by Geyser 1000 W "=1000*1=1kWh"

Energy consumed by Ac-1500 W "=1500*4=6kWh"

Total energy consumed= "0.32 + 0.48+1+6= 7.8 kWh"

The energy bill for the month of March "=Rs 5.25\/-*7.8kWh=Rs 40.95\/-"